eelyaK11:

A NASA spacecraft once landed on the surface of the asteroid named 433 Eros.This asteroid is near Earth and has an average orbital speed of 24.36 kilometers per second. What is its speed in meters per second? 0.02436 meters per second 0.2436 meters per hour 2,460 meters per hour 24,360 meters per second

eelyaK11:

is it b?

MxxnLight:

A NASA spacecraft once landed on the surface of the asteroid named 433 Eros. This asteroid is near Earth and has an average orbital speed of 24.36 kilometers per second.

MxxnLight:

Hmmm.

eelyaK11:

a?

eelyaK11:

or b?

MxxnLight:

I would go with..

MxxnLight:

24,360 meters per second I may be wrong.

eelyaK11:

u right

MxxnLight:

i- ok.

jimthompson5910:

This is what your steps could look like \(\large 24.36 \text{ km/s} = \frac{24.36 \text{ km}}{1 \text{ sec}}\times\frac{1000 \text{ m}}{1 \text{ km}}\) Note how the "km" units cancel \(\large 24.36 \text{ km/s} = \frac{24.36 \cancel{\text{ km}}}{1 \text{ sec}}\times\frac{1000 \text{ m}}{1\cancel{\text{ km}}}\) Leaving us with this \(\large 24.36 \text{ km/s} = (24.36*1000) \text{ m/s}\) \(\large 24.36 \text{ km/s} = 24,360 \text{ m/s}\)